Calculus II, Fall 2014 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 11

1. Section 14.10: Problems 3, 7, 9, 12.

2. Section 15.1: Problems 7, 19.

3. Section 15.2: Problems 11, 19, 27, 29, 35, 43, 47, 69.

4. Section 15.3: Problems 13, 17.